# Astrobites Collaboration Improving Undergraduate Astronomy Students’ Skills with Research Literature via Accessible Summaries: A Case Study with Astrobites-based Lesson Plans

### v2

@@ V Conclusions and Recommendations
Overall, our results show that the Astrobites lesson plans were largely successful at their main goal–increasing student comfort and ability with research literature–plus other beneficial effects, such as increasing student knowledge about the field of astronomy and the process of research. In particular, our reading-intensive intervention (Lesson Plan 1) significantly improved student ability and confidence with parsing jargon and understanding the main takeaways of an article. Our results, therefore, also show strong evidence for the positive benefits of reading comprehension assignments featuring accessible summaries of current literature more generally; as such, we identify a recommendation for instructors to use such assignments and a need for further research work in this area to fully maximize the benefits and understand the mechanisms at play in student learning.
Open-ended responses from participants also indicate that the accessible summary was a key component of the lesson plans’ success in increasing student confidence with literature; scaffolding steps have long been shown to improve student outcomes [52], and our work indicates a similar phenomenon for the task of reading complex scientific articles. Scaffolding scientific research articles can often be difficult, as the available alternatives to cover the content (namely science journalism) often reduce the complexity to the point of being an entirely different, non-technical genre that likely cannot target the same skills of reading technical literature. Astrobites and similar efforts represent the creation of a third writing type in the field, the accessible summary, intended for beginner technical audiences such as the students in this study population. Our results indicate that this type of writing is strongly beneficial for students as they enter the realm of scientific research.
@@
As the sample in this paper focused on reading research papers, however, there is still significant work to be done to explore how writing lessons (e.g. Lesson Plan 3) impact students in STEM classrooms; we envision this as complimentary work to the reading comprehension interventions we studied, especially as reading is the first step to writing [53]. Additionally, as mentioned previously, it is yet unclear how students’ gains will change with repeated exposure to these lessons, e.g., if the majority of change is in the first exposure or if they grow substantially with practice; future work will need to test these “dosing effects” to determine the optimal usage of these lessons.
@@
Accessible summaries show significant promise for bridging the gap between students’ entry points and the difficult task of reading scientific literature. Additionally, it is important to note that instructors were appreciative of the existence of pre-made resources to use=, as opposed to starting from scratch to incorporate reading in their courses; future efforts in this area should similarly work to provide instructors with usable lessons to encourage uptake of reading/writing-focused activities in astronomy courses. We strongly encourage further work on integrating reading comprehension exercises and research literature into the classroom, especially through accessible summaries as in the Astrobites lesson plan models shown herein to positively benefit students.
@@ Acknowledgements.
B.L.L. acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. 2021-25 DGE-2034835. A.R.W. acknowledges support from the Virginia Space Grant Consortium and the National Science Foundation Graduate Research Fellowship Program under grant No. 1842490. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s)and do not necessarily reflect the views of the National Science Foundation. Support for C.J.L. was provided by NASA through the NASA Hubble Fellowship grant No. HST-HF2-51535.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. RRL is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the form of an Emmy Noether Research Group – Project-ID 445674056 (SA4064/1-1, PI Sander). This work was partially funded by the American Astronomical Society Education and Professional Development (EPD) mini-grant program. This study was conducted with a UCLA IRB Exemption Certificate IRB#22-001473 and UVA IRB Approval Certificate IRB#5514. B. Lewis would like to thank the UCLA CIRTL TAR program for their guidance and mentorship, especially her CIRTL co-authors on this paper, Rachel Kennison for serving as faculty sponsor for UCLA IRB, and Graham Read. Thank you to Nicholas Young for valuable feedback on this manuscript, Jessie Thwaites and Haley Wahl for participating in talk-throughs as part of the survey design process, to Luna Zagorac for proofreading, to Nathan Sanders for his unwavering support of Astrobites’ education initiatives, to Huei Sears for her support of grant proposals to fund this work, and to the rest of the Astrobites collaboration for their support of this project.References
@@
- Barthelemy et al. [2022] R. S. Barthelemy, M. Swirtz, S. Garmon, E. H. Simmons, K. Reeves, M. L. Falk, W. Deconinck, E. A. Long, and T. J. Atherton, LGBT + physicists: Harassment, persistence, and uneven support, Physical Review Physics Education Research 18, 010124 (2022).
- James et al. [2020] W. James, C. Bustamante, K. Lamons, E. Scanlon, and J. J. Chini, Disabling barriers experienced by students with disabilities in postsecondary introductory physics, Phys. Rev. Phys. Educ. Res. 16, 020111 (2020).
- Kricorian, Katherine and Seu, Michelle, and Lopez, Daniel and Ureta, Elsie and Equils, Ozlem [2020] Kricorian, Katherine and Seu, Michelle, and Lopez, Daniel and Ureta, Elsie and Equils, Ozlem, Factors influencing participation of underrepresented students in STEM fields: matched mentors and mindsets, International Journal of STEM Education 7, 7 (2020).
@@
- Lewis et al. [2022] B. L. Lewis, K. Supriya, G. H. Read, K. L. I. Dixie, R. Kennison, and A. R. Friscia, Effects of popular science writing instruction on general education student attitudes towards science: A case study in astronomy, arXiv preprint arXiv:2209.02852 10.48550/arXiv.2209.02852 (2022).
@@
- Note [1] astrobites.org.
- Note [2] Related majors include physics, astrophysics, biology, math, geology, engineering, chemistry, computer science, planetary science, and undeclared majors.
@@
- Cohen [1988] J. Cohen, Statistical Power Analysis for the Behavioral Sciences (Lawrence Erlbaum Associates, 1988).
@@
- Note [3] For quantitative questions 2 and 5, low Likert scores (disagreement) represent greater confidence. This is the opposite of the majority of the questions, for which higher Likert scores indicate greater confidence.
@@
- Yardley [2015] L. Yardley, Demonstrating validity in qualitative psychology, Qualitative psychology: A practical guide to research methods 3, 257 (2015).
@@
- De Winter et al. [2016] J. C. F. De Winter, S. D. Godwin, and J. Potter, Comparing the pearson and spearman correlation coefficients across distributions and sample sizes: A tutorial using simulations and, Psychological Methods , 273 (2016).
@@ Appendix A About Astrobites
The Astrobites collaboration is a group of volunteer graduate students in astronomy and related fields who publish daily blog-style posts summarizing current astronomy research papers. Astrobites also publishes guides to specific topics and other resources relevant to career development and equity/inclusion. While Astrobites content is primarily targeted towards undergraduate astronomy majors, its audience has grown to include everyone from hobbyists to emeritus professors since our founding in 2011.
A central focus of Astrobites’ mission is to further efforts for inclusion and equity in astronomy. We impact accessibility by addressing the inaccessibility and technical language of research articles through summaries, and our host of “guides” to different aspects of astronomy and academic professional development. Astrobites articles are also an excellent way for students to explore differences in audience between various forms of science writing, and for students–both undergraduates reading the articles and graduate students serving as regular writers–to practice writing skills, which are increasingly critical for becoming a successful scientist. Astrobites has also taken an active role in incorporating its pedagogical content, wide topical coverage, and scaffolded contextual resources into formal education [31, 32].
@@ Appendix B Full Assessment Text
Appendix C Further Information on Lesson Plans
@@
Astrobites currently maintains four lesson plans based on our prior published work [32]. These lesson plans provide four pathways to engage students with Astrobites materials—guided reading assignments, student research projects, a student writing assignment, and student presentations—and are available in this linked document. For each lesson plan, we provide an overview of the learning objectives, an outline of how the plan can be adapted for different course levels, sample handouts for presenting the assignment to students, and a potential grading rubric.
@@
The first lesson plan (Lesson Plan 1: Periodic Reading Assignments), which features a series of guided reading assignments, is designed to grow students’ reading comprehension and strengthen their conceptual understanding of course material in the broader context of active research. Students build familiarity and confidence in their ability to monitor and keep pace with recent developments in the astronomy literature by reading a series of Astrobites articles and testing their understanding through directed questions. This lesson can be easily adapted to different course levels by varying the frequency of assignments and altering the style of the associated questions (e.g., asking more directed questions for lower-division undergraduates compared to open-ended questions for graduate students).
@@
In the second lesson plan (Lesson Plan 2: Student Research Project), students select a topic and prepare a paper or presentation using Astrobites articles to guide their research. Students synthesize concepts and interpret figures and data from different articles for their report. Students build confidence in developing techniques to tackle new concepts by using Astrobites articles to identify technical terminology (jargon) with which they are unfamiliar. Students then define jargon by creating a glossary. This lesson plan is meant to teach students how to engage with full research articles in the future.
@@
For the third lesson plan (Lesson Plan 3: Student Writing Assignment), students read one or more papers and write an Astrobites-like summary. Much as with the second lesson plan, students practice synthesizing information from multiple sources. However, they additionally develop valuable science communication and composition skills by compiling information into a bite-sized article. Additionally, students actively practice reading the scientific literature directly. By scaffolding the assignment—e.g., having students present drafts and outlines along the way—instructors can help students settle into the reading and writing process more smoothly.
@@
Finally, in the fourth lesson plan (Lesson Plan 4: Student Presentation), students deliver an in-class presentation featuring a profile of an astronomer who has been interviewed by Astrobites and a summary of a paper by the interviewee. While also incorporating the reading comprehension and communication skills that appear in the other lesson plans, the focus of this lesson plan is to support students’ career preparation and develop a sense of belonging in the field. By sharing presentations describing the career trajectory and advice of successful astronomers hailing from diverse backgrounds, students are able to see that astronomy is for everyone, while also identifying and discussing important steps on the way to becoming a professional astronomer. For example, one study of the importance of role models for career preparation and aspiration of young Black British people is presented in Archer et al.[54] and summarized in Astrobites in 2020.
Appendix D Additional Response Quotes from Students and Educators
Student comments on the usefulness of the Astrobites lesson plans:
“The activity helped me to understand the findings of a research paper at my level of knowledge. It wasn’t too technical, and wasn’t too ‘dumbed-down’ either.”
“While the stuff we learn in class is important, I felt like this activity took we learned in class and made us do something ‘practical’ with it.”
“I learned more about other astronomers and their contributions (no matter how big or small) to astronomy.”
“I found the length of the Astrobites less daunting to read about and made the experience more enjoyable.”
Instructor comments on the utility of the Astrobites lesson plans:
@@ Appendix D Additional Response Quotes from Students and Educators
“[Astrobites] enabled the possibility to connect to the paper, because the paper was so inaccessible for the current know-how of the students.”
@@
Appendix E Further information on quantitative analysis
This appendix includes histograms (Figure 6) and a table of statistical summaries (Table 5) for all of the Likert response questions individually, supplementary to Section III.1.
Q | CI95% | p | cohen-d | d interp | BF10 | BF10 interp | power |
1 | [0.25 0.75] | 1.27 | 0.456 | small | 157 | extreme for H1 | 0.983 |
2 | [-0.39 0.17] | 0.439 | 0.0702 | very small | 0.163 | moderate for H0 | 0.0962 |
3 | [-0.08 0.39] | 0.184 | 0.156 | very small | 0.288 | moderate for H0 | 0.286 |
4 | [0.09 0.55] | 7.82 | 0.259 | small | 3.83 | moderate for H1 | 0.638 |
5 | [-0.58 0.15] | 0.236 | 0.131 | very small | 0.242 | moderate for H0 | 0.217 |
6 | [0.3 0.79] | 2.56 | 0.428 | small | 698 | extreme for H1 | 0.969 |
7 | [-0.07 0.46] | 0.149 | 0.175 | very small | 0.337 | anecdotal for H0 | 0.348 |
8 | [-0.17 0.34] | 0.502 | 0.0688 | very small | 0.152 | moderate for H0 | 0.0944 |
9 | [-0.05 0.44] | 0.114 | 0.175 | very small | 0.412 | anecdotal for H0 | 0.347 |
10 | [0.02 0.57] | 0.0373 | 0.232 | small | 1.01 | anecdotal for H1 | 0.547 |
11 | [0.03 0.53] | 0.0267 | 0.245 | small | 1.33 | anecdotal for H1 | 0.591 |
12 | [0.05 0.53] | 0.0177 | 0.290 | small | 1.89 | anecdotal for H1 | 0.737 |
13 | [-0.03 0.39] | 0.0874 | 0.179 | very small | 0.506 | anecdotal for H0 | 0.361 |
14 | [-0.11 0.36] | 0.305 | 0.107 | very small | 0.203 | moderate for H0 | 0.160 |
15 | [-0.05 0.47] | 0.113 | 0.150 | very small | 0.413 | anecdotal for H0 | 0.267 |
16 | [-0.08 0.44] | 0.163 | 0.156 | very small | 0.314 | moderate for H0 | 0.287 |
17 | [-0.23 0.32] | 0.726 | 0.0359 | very small | 0.129 | moderate for H0 | 0.0619 |
18 | [-0.03 0.47] | 0.0803 | 0.160 | very small | 0.540 | anecdotal for H0 | 0.298 |
19 | [-0.05 0.46] | 0.110 | 0.158 | very small | 0.422 | anecdotal for H0 | 0.294 |
20 | [-0.13 0.32] | 0.386 | 0.0749 | very small | 0.176 | moderate for H0 | 0.103 |
21 | [-0.19 0.31] | 0.625 | 0.0413 | very small | 0.137 | moderate for H0 | 0.0658 |
22 | [-0.09 0.31] | 0.274 | 0.0987 | very small | 0.218 | moderate for H0 | 0.143 |
23 | [-0.18 0.23] | 0.813 | 0.0198 | very small | 0.125 | moderate for H0 | 0.0536 |
24 | [0.22 0.71] | 3.51 | 0.324 | small | 62.1 | very strong for H1 | 0.827 |
25 | [-0.11 0.33] | 0.332 | 0.0889 | very small | 0.193 | moderate for H0 | 0.125 |
26 | [-0.22 0.22] | 1.00 | 0.00 | very small | 0.122 | moderate for H0 | 0.05 |
27 | [-0.07 0.48] | 0.135 | 0.151 | very small | 0.361 | anecdotal for H0 | 0.273 |
28 | [0.18 0.72] | 1.45 | 0.347 | small | 17.2 | strong for H1 | 0.873 |
29 | [0. 0.58] | 0.0479 | 0.223 | small | 0.821 | anecdotal for H0 | 0.513 |
30 | [-0.16 0.33] | 0.489 | 0.0695 | very small | 0.154 | moderate for H0 | 0.0954 |
31 | [-0.11 0.35] | 0.295 | 0.104 | very small | 0.208 | moderate for H0 | 0.153 |
32 | [-0.18 0.31] | 0.622 | 0.0418 | very small | 0.137 | moderate for H0 | 0.0662 |
33 | [-0.1 0.37] | 0.262 | 0.0902 | very small | 0.225 | moderate for H0 | 0.127 |
34 | [-0.17 0.37] | 0.472 | 0.0634 | very small | 0.157 | moderate for H0 | 0.0875 |
35 | [0.08 0.68] | 0.0142 | 0.237 | small | 2.28 | anecdotal for H1 | 0.562 |
Appendix F Consistency Analylsis
When performing a mixed methods study such as this, it is good practice to compare results obtained through multiple questions covering the same themes. One way to conduct such a comparison involves assigning numerical scores to qualitative questions [55, 56]. Here, we identified three qualitative questions (Questions 36, 38, and 39 in Appendix B) for which there are Likert scale questions that address similar themes. We determined which qualitative response codes denoted positive, neutral, and negative feelings toward the three questions, and assigned each student a qualitative “score” for that question based on the prevalence of each type of code in their response. The related Likert questions and positive, neutral, and negative codes considered for each of the three qualitative questions used in this comparison are listed in Table 6.
For each student response to each of these three questions, we took the average of related Likert scale questions to find their “Likert score” (quantitative). To assign a student’s “written score” (qualitative) on a topic, we assigned to every positive code applied to their response to that question, to every negative code, and to every neutral code. By adding these values, we found a total score — if this total score was positive (negative), we assigned the student a written score of () for that question; if the total score was , we left it as . For each of the three qualitative questions selected, we compared each student’s Likert and written scores, as well as the score change for each student from pre- to post-survey.
Note that not every Likert question was used in this analysis, and several Likert questions were considered in relation to more than one qualitative question. We only considered codes applied by a majority of raters, as in the general qualitative analysis, and we considered the number of raters who applied a given code (e.g., if three scorers applied a positive code to a response but five applied a negative code, the response received a and was scaled to a written score of ). We experimented with taking the average value of the codes (e.g., the former example would yield a score of ), though we found that the correlation between Likert and written scores was not significantly impacted by this change in methodology. We proceeded to use the integer scores () in recognition of the inherent subjectivity of assigning numerical scores to qualitative responses based on response codes.
Open-Ended Question | Codes considered | Related Likert questions |
36 (Papers) | Int, U, E, | 1, 2, 3, 4, 5, 7, 8, 9, 11, 28 |
Co, W, En | ||
38 (Capable) | Y, N, M | 14, 18, 25, 26, 27, 28, 29, 30, 31, 32 |
39 (Belonging) | Y, N, M | 32, 33, 34, 35 |
Written response scores tend to track Likert scores for each of the questions (Figure 8); however, the correlation between the two scores, as computed with the Spearman correlation coefficient, is smaller in the post-survey than in the pre-survey. This is true for all three questions for which we performed consistency analysis. Many students show no change in written score from pre- to post-survey (Figure 9), though many students, for each collection of Likert scale questions, exhibit a positive change after the activities.
Appendix G Correlation Analysis
Quantitative survey questions were used to study second-order effects, how improvements or reductions in an ability to parse information correlate to changes in an ability to apply that knowledge. We divided the quantitative questions into four categories: one that represented understanding of information (“Understanding”), one that represented synthesizing that information into more abstract concepts (“Synthesizing”), one that represented discussing the material with others (“Communication”), and one that represented how students viewed themselves in the context of the field of astronomy and the degree to which they identified as being part of the astronomy community (“Self-Assessment”).
Details of the makeup of each category are listed in Table 7. Note that not all questions were used and that these categories are distinct from those used in the assessment design (the “conceptual catgeories”) as detailed in Section II and summarized in Table 1.
Category | Question numbers in category |
Understanding | 1, 3, 4, 7, 8, 9, 10, 13, 18 |
Synthesizing | 11, 12, 13, 14, 15, 16, 17 |
Communicating | 6, 19, 20, 21, 22, 23, 24 |
Self-assessment | 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35 |
For each student, we computed their initial score in each category by averaging their responses to questions in the category from the pre-lesson survey. We computed the student’s final score in the category the same way, using responses from the post-lesson survey. We then quantified changes in each category by taking the difference of the two scores. The correlations between category improvement or reduction were described quantitatively by the Spearman correlation coefficient. We choose to use the Spearman coefficient and not the Pearson correlation coefficient, because the Likert scale is ordinal, meaning that any relationships may not be linear, and it is possible that there are outliers due to, e.g., students choosing to not take the survey seriously. This combination makes the Spearman coefficient the better choice [57].
In all categories, the majority of students showed no substantial net change, but students who did exhibit a change in skill largely improved. Figure 5 is a Sankey diagram of the “understanding” category showing the evolution of individual students, broken down by pre-activity score in that area. Sankey diagrams for the remaining three categories show similar trends. Students with low pre-activity scores generally exhibited slight growth; students with medium pre-activity scores generally exhibited no substantial change or slight growth; and students with high pre-activity scores generally exhibited no substantial change or slight reduction.
We found mild positive correlations between all four skill categories, with the strongest being between “Understanding” and “Synthesizing” (Spearman coefficient ). No correlations were negative; for individual students, an increase in one area seldom corresponded with a decrease in another. Table 8 displays the Spearman coefficients for each combination of skill categories.
Category | Understanding | Synthesizing | Communicating | Self-assessment |
Understanding | N/A | 0.660 | 0.487 | 0.540 |
Synthesizing | 0.660 | N/A | 0.511 | 0.410 |
Communicating | 0.487 | 0.511 | N/A | 0.416 |
Self-assessment | 0.540 | 0.410 | 0.416 | N/A |
We hypothesized that changes in self-assessment scores may be a function of changes in the three learning categories (understanding, communicating, and synthesizing). To test this, we performed multiple linear regression to determine the coeffients for the following equation:
The variables showed moderate dependency on each other, displaying Variance Inflation Factors of 2-3. This is considered moderate dependency, and as such we perform standard multiple linear regression. Our results are shown in the following table:
Coefficient | Value | Error | P-value |
0.017 | 0.069 | 0.803 | |
0.473 | 0.145 | 0.002 | |
0.505 | 0.129 | 0.001 | |
-0.047 | 0.143 | 0.743 |
We find the relationships between self-assessment and the other three learning categories to be statistically significant, primarily the relationships between self-assessment, understanding, and communicating.

### v3

@@ VII Conclusions & Recommendations
Overall, our results show that the accessible summary reading assignment intervention was largely successful at its main goal – increasing student comfort and perceived ability with research literature – and yielded other beneficial effects, such as increasing student knowledge about the field of astronomy and the process of research. This intervention significantly improved student ability and confidence with parsing jargon and understanding the main takeaways of a research summary article, which may translate to improved ability with primary research literature. Our results, therefore, show evidence for the promise of positive benefits of reading comprehension assignments featuring accessible summaries of current literature more generally, although it is not yet known how this intervention compares to other reading comprehension exercises. As such, we identify a recommendation for instructors to use such assignments and a need for further research work in this area to fully maximize the benefits and understand the mechanisms at play in student learning.
@@
It is important to note that this study focused on a relatively small sample of students, and there is likely a bias in our sample due to the self-selecting nature of instructors who chose to participate (i.e. motivated educators seeking out new pedagogical strategies). Further work will be needed to make truly generalizable conclusions about the effects of reading comprehension interventions on student ability with research literature and other outcomes. As mentioned previously, it is also yet unclear how students’ gains will change with repeated exposure to these lessons, e.g., if the majority of change is in the first exposure or if they grow substantially with practice. Future work will need to test these “dosing effects” to determine the optimal usage of these lessons. The sample in this paper focused on the skills necessary for the task of reading research papers, however, there is also still significant work to be done to explore how writing lessons impact students in STEM classrooms. We envision this as complimentary work to the reading comprehension interventions we studied, especially as reading is the first step to writing [63].
Open-ended responses from participants also indicate that the accessible summary was a key component of the assignment’s success in increasing student confidence with literature; scaffolding steps have long been shown to improve student outcomes [64], and our work indicates a similar phenomenon for the task of reading complex scientific articles. Scaffolding students’ comprehension of scientific research articles can often be difficult, as the available alternatives to cover the content (namely science journalism) often reduce the complexity to the point of being an entirely different, non-technical genre that likely cannot target the same skills of reading technical literature. Astrobites and similar efforts represent the creation of a third writing type in the field, the accessible summary, intended for beginner technical audiences such as the students in this study population. Our results indicate that this type of writing may be beneficial for students as they enter the realm of scientific research.
@@
In conclusion, accessible summaries show promise for bridging the gap between students’ entry points and the difficult task of reading scientific literature, although significant further study is needed to expand upon this exploratory case study. Additionally, it is important to note that instructors were appreciative of the existence of pre-made resources to use, as opposed to starting from scratch to incorporate reading in their courses. Future efforts in this area should similarly work to provide instructors with usable activities and templates to encourage uptake of reading/writing-focused activities in astronomy courses. We strongly encourage further work on integrating reading comprehension exercises and research literature into the classroom, especially through accessible summaries as in this exploratory case study.
@@ Acknowledgements.
B.L.L. acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. 2021-25 DGE-2034835 and the National Science Foundation Astronomy & Astrophysics Postdoctoral Fellowship under Award No. 2401654. A.R.W. acknowledges support from the Virginia Space Grant Consortium and the National Science Foundation Graduate Research Fellowship Program under grant No. 1842490. S.G. acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. 2233001. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s)and do not necessarily reflect the views of the National Science Foundation. Support for C.J.L. was provided by NASA through the NASA Hubble Fellowship grant No. HST-HF2-51535.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. RRL is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the form of an Emmy Noether Research Group – Project-ID 445674056 (SA4064/1-1, PI Sander). This work was partially funded by the American Astronomical Society Education and Professional Development (EPD) mini-grant program. This study was conducted with a UCLA IRB Exemption Certificate IRB#22-001473 and UVA IRB Approval Certificate IRB#5514. B. Lewis would like to thank the UCLA CIRTL TAR program for their guidance and mentorship, especially her CIRTL co-authors on this paper, Rachel Kennison for serving as faculty sponsor for UCLA IRB, and Graham Read. Thank you to Nicholas Young for valuable feedback on this manuscript, Jessie Thwaites and Haley Wahl for participating in talk-throughs as part of the survey design process, to Luna Zagorac for proofreading, to Nathan Sanders for his unwavering support of Astrobites’ education initiatives, to Huei Sears for her support of grant proposals to fund this work, to Alexandra Masegian for her assistance in coding data, and to the rest of the Astrobites collaboration for their support of this project.References
@@
- Lewis et al. [2023] B. L. Lewis, K. Supriya, G. H. Read, K. L. I. Dixie, R. Kennison, and A. R. Friscia, Effects of popular science writing instruction on general education student attitudes towards science: A case study in astronomy, Astronomy Education Journal 10.32374/AEJ.2023.3.1.049ra (2023).
@@
- Baram-Tsabari and Yarden [2011] A. Baram-Tsabari and A. Yarden, Quantifying the gender gap in science interests, International Journal of Science and Mathematics Education 9, 523 (2011).
@@
- Pepper et al. [2018] D. Pepper, J. Hodgen, K. Lamesoo, P. Kõiv, and J. Tolboom, Think aloud: using cognitive interviewing to validate the pisa assessment of student self-efficacy in mathematics, International Journal of Research & Method in Education 41, 3 (2018).
- Wolcott and Lobczowski [2021] M. D. Wolcott and N. G. Lobczowski, Using cognitive interviews and think-aloud protocols to understand thought processes, Currents in Pharmacy Teaching and Learning 13, 181 (2021).
- Formanek et al. [2019] M. Formanek, S. Buxner, C. Impey, and M. Wenger, Relationship between learners’ motivation and course engagement in an astronomy massive open online course, Phys. Rev. Phys. Educ. Res. 15, 020140 (2019).
- Zou et al. [2024] Y. Zou, X. Xue, L. Jin, X. Huang, and Y. Li, Assessment of conceptual understanding in student learning of evaporation, Phys. Rev. Phys. Educ. Res. 20, 020107 (2024).
- Ding et al. [2024] Y. Ding, G. Zhu, Q. Bian, and L. Bao, Analysis of students’ conceptual change in learning newton’s third law with an integrated framework of model analysis and knowledge integration, Phys. Rev. Phys. Educ. Res. 20, 020141 (2024).
- Cronbach [1951] L. J. Cronbach, Coefficient alpha and the internal structure of tests, Psychometrika 16, 297 (1951).
- AIP [2024a] (2024a).
- AIP [2024b] (2024b).
- Zhai et al. [2010] F. Zhai, C. C. Raver, S. M. Jones, C. P. Li-Grining, E. Pressler, and Q. Gao, Dosage effects on school readiness: Evidence from a randomized classroom-based intervention, Social Service Review 84, 615 (2010).
@@
- Lee and Wagenmakers [2014] M. D. Lee and E.-J. Wagenmakers, Bayesian cognitive modeling: A practical course (Cambridge university press, 2014).
@@
- Gisev et al. [2013] N. Gisev, J. S. Bell, and T. F. Chen, Interrater agreement and interrater reliability: Key concepts, approaches, and applications, Research in Social and Administrative Pharmacy 9, 330 (2013).
- Grayson and Rust [2001] K. Grayson and R. Rust, Interrater reliability, Journal of Consumer Psychology 10, 71 (2001).
- Campbell et al. [2013] J. L. Campbell, C. Quincy, J. Osserman, and O. K. Pedersen, Coding in-depth semistructured interviews: Problems of unitization and intercoder reliability and agreement, Sociological Methods & Research 42, 294 (2013), https://doi.org/10.1177/0049124113500475 .
- Stemler [2004] S. E. Stemler, A comparison of consensus, consistency, and measurement approaches to estimating interrater reliability, Practical Assessment, Research and Evaluation 9, 4 (2004).
- Kurasaki [2000] K. S. Kurasaki, Intercoder reliability for validating conclusions drawn from open-ended interview data, Field Methods 12, 179 (2000), https://doi.org/10.1177/1525822X0001200301 .
- Holdford [2008] D. Holdford, Content analysis methods for conducting research in social and administrative pharmacy, Research in Social and Administrative Pharmacy 4, 173 (2008).
@@ Appendix A About Astrobites
The Astrobites collaboration (or simply Astrobites) is a group of volunteer graduate students in astronomy and related fields who publish daily blog-style posts summarizing current astronomy research papers. Astrobites also publishes guides to specific topics and other resources relevant to career development and equity/inclusion. While the content from the Astrobites website is primarily targeted towards undergraduate astronomy majors, its audience has grown to include everyone from hobbyists to emeritus professors since its founding in 2011.
A central focus of Astrobites’ mission is to further efforts for inclusion and equity in astronomy. The organization aims to address the inaccessibility and technical language of research articles through summaries and “guides” to different aspects of astronomy and academic professional development. Articles featured on Astrobites are also an excellent way for students to explore differences in audience between various forms of science writing, and for students – both undergraduates reading the articles and graduate students serving as regular writers – to practice writing skills. These skills are increasingly critical for becoming a successful scientist. Astrobites has also taken an active role in incorporating its pedagogical content, wide topical coverage, and scaffolded contextual resources into formal education [29, 30].
@@ Appendix B Full Assessment Text
Appendix C Further Information on Astrobites Assignments
@@
Astrobites currently maintains four assignment templates based on prior published work [30]. These templates provide four pathways to engage students with material from Astrobites – guided reading assignments, student research projects, a student writing assignment, and student presentations – and are available in this linked document. For each assignment, an overview of the learning objectives, an outline of how the plan can be adapted for different course levels, sample handouts for presenting the assignment to students, and a potential grading rubric are provided for instructors.
@@
The first assignment (Periodic Reading Assignments) is designed to grow students’ reading comprehension while strengthening their conceptual understanding of course material in the broader context of active research. Students build familiarity and confidence in their ability to monitor and keep pace with recent developments in the astronomy literature by reading a series of Astrobites articles and testing their understanding through directed questions. This assignment can be easily adapted to different course levels by varying the frequency of assignments and altering the style of the associated questions (e.g., asking more directed questions for lower-division undergraduates compared to open-ended questions for graduate students).
@@
In the second assignment (Student Research Project), students select a topic and prepare a paper or presentation using articles from the Astrobites website to guide their research in a modified literature review. Students synthesize concepts and interpret figures and data from different articles for their report. Students build confidence in developing techniques to tackle new concepts by using Astrobites articles to identify technical terminology (jargon) with which they are unfamiliar. Students then define jargon by creating a glossary. This assignment is meant to teach students how to engage with full research articles in the future.
@@
For the third assignment (Student Writing Assignment), students read one or more papers and write an Astrobites-like summary. Much as with the second assignment, students practice synthesizing information from multiple sources. However, they additionally develop valuable science communication and composition skills by compiling information into a bite-sized article. Additionally, students actively practice reading the scientific literature directly. By breaking the assignment into smaller chunks – e.g., having students present drafts and outlines along the way – instructors can help students settle into the reading and writing process more smoothly.
@@
Finally, in the fourth assignment (Student Presentation), students deliver an in-class presentation featuring a profile of an astronomer who has been interviewed by people from the Astrobites collaboration and a summary of a paper by the interviewee. While also incorporating the reading comprehension and communication skills that appear in the other assignments, the focus of this assignment is to support students’ career preparation and develop a sense of belonging in the field. By sharing presentations describing the career trajectory and advice of successful astronomers hailing from diverse backgrounds, students are able to see that astronomy is for everyone, while also identifying and discussing important steps on the way to becoming a professional astronomer. For example, one study of the importance of role models for career preparation and aspiration of young Black British people is presented in ARCHER et al. [65] and was summarized in Astrobites in 2020.
Appendix D Additional Response Quotes from Educators
Instructor comments on the utility of the assignment templates featuring accessible summaries:
@@ Appendix D Additional Response Quotes from Educators
“[Astrobites/accessible summaries] enabled the possibility to connect to the paper, because the paper was so inaccessible for the current know-how of the students.”